OpenStudy (anonymous):

Help! How do I rationalize the denominator in this problem : 3rd root of 1/6 , so that the result is in simplest form?

OpenStudy (anonymous):

= 6 ^(2/3) / 6

OpenStudy (anonymous):

How do you get 6^(2/3) out of \[\sqrt[3]{1/6}\]

OpenStudy (anonymous):

I get 3rd root of 6/6

OpenStudy (accessdenied):

The idea here is to break it apart like this since \(\sqrt{\frac{a}{b}} = \frac{\sqrt{a}}{\sqrt{b}} \). \[ \sqrt[3]{\frac{1}{6}} = \frac{\sqrt[3]{1}}{\sqrt[3]{6}} \] To rationalize this, we need to multiply the numerator and the denominator by a factor that cancels the cube root in the denominator. We know that the number on the inside has to be a \(\textbf{perfect cube}\) to simplify, so we have to think about what we can multiply to \(\sqrt[3]{6}\) to create that perfect cube, and then multiply it to both the numerator and the denominator..

OpenStudy (anonymous):

would I multiply the numerator and denominator by \[\sqrt[3]{6}?\]

OpenStudy (anonymous):

that was my first thought when I wrote the problem down, to multiply by \[\sqrt[3]{6}\]

OpenStudy (accessdenied):

Well, what happens then, is that we get this: \[ \frac{\sqrt[3]{6}}{\sqrt[3]{6^2}} = \frac{\sqrt[3]{6}}{\color{green}{\sqrt[3]{36}}}\] But is 36 a perfect cube that we can take out? Or maybe we need to multiply by another \(\sqrt[3]{6}\) to make it a perfect cube? :)

OpenStudy (accessdenied):

If we multiply by another cube root of 6 to both numerator and denominator, we create a \(\large \sqrt[3]{6^3}\) in the denominator, so the cube root and the cube cancel to leave 6. \[ \frac{\sqrt[3]{36}}{6} \] That numerator is also 6^2/3, as previously mentioned.

OpenStudy (anonymous):

oh, when I said earlier, "would I multiply the numerator & denominator by \[\sqrt[3]{6}\] " ...isn't that what you just said now? "to multiply by another cube root of 6 to both numerator and denominator"?

OpenStudy (accessdenied):

The idea is to make a perfect cube under the cube root with the denominator. When we multiplied by the first cubed root of 6, we got cuberoot(36) in the denominator, which isn't a perfect cube. We had to instead multiply by another cuberoot(6) to make it a perfect cube. (We multiply 6*6*6, 6^3, so that's a perfect cube). Effectively, we multiplied the numerator and the denominator by \(\sqrt[3]{6^2}\) , or 6^(2/3), since we had 6^(1/3) already. \[ \frac{1}{\sqrt[3]{6}} \times \frac{\sqrt[3]{6^2}}{\sqrt[3]{6^2}} \]

OpenStudy (accessdenied):

Unlike square roots, which you just need one other square root to make it 'complete': \(\sqrt{2} \times \sqrt{2} = 2\) Cube roots need three: \(\sqrt[3]{2} \times \sqrt[3]{2} \times \sqrt[3]{2} = 2\)

OpenStudy (anonymous):

ok, the denominator would be 6, and the numerator would be \[\sqrt[3]{36}\]?

OpenStudy (accessdenied):

Yeah. That's correct.

OpenStudy (anonymous):

I think I see that now, if the denominator is, for example, \[\sqrt[3]{2}\], you would multiply numerator and denominator by \[\sqrt[3]{2^{2}}\] ?

OpenStudy (accessdenied):

Yeah. We complete the cube under the cube root and it pretty much just cancels with the root to get the whole number.

OpenStudy (anonymous):

What if the denominator is for example, \[\sqrt[5]{8}\] would you multiply numerator and denominator by \[\sqrt[5]{8^{4}}\] ?

OpenStudy (accessdenied):

Yep. Or, if you think about it, \(8=2^3\), so we could also multiply by \(\sqrt[5]{2^2}\) since that makes \(\sqrt[5]{2^5}\). If you do it your way, usually you just have to do some extra simplifying and you'll get the simplest form.

OpenStudy (accessdenied):

Little tricks like that come up every so often when you have some perfect squares, cubes, etc. as factors.

OpenStudy (anonymous):

The key is we want exponent is 1 => 1 - 1/5 = 4/5

OpenStudy (anonymous):

By fraction rule, if we multiply/ divide => it need to be done on BOTH upper and lower!

OpenStudy (anonymous):

@AccessDenied you always deserve multi medals more than I'm ;)

OpenStudy (anonymous):

*deserved

OpenStudy (anonymous):

Thank you for your help, AccessDenied & Chlorophyll!!...both of you have made this type of problem easier to understand1

OpenStudy (anonymous):

I just copy cat the idea after @AccessDenied :P

OpenStudy (anonymous):

when something is raised to a fractional exponent, it makes the problem seem difficult, at first! ...thanks again for your help!

OpenStudy (accessdenied):

You're welcome! The rational exponents, I find, are usually easier to represent some of the properties like that previous cancelling the square and sixth root. It's just harder to imagine a power as a fraction since it usually means "take this multiplied to itself # times," but as a fraction, its like "take this multiplied to itself but only half-way". The fraction just indicates that we don't have the 'whole' thing there.